OpenStudy (anonymous):

solve system of equation: x+y=-1 y-z=4 2x+y+5z=-38 i need step by step help please?

OpenStudy (anonymous):

step1; multiply equation 1 with -1

OpenStudy (anonymous):

step2; add equation 1 and equation 2

OpenStudy (anonymous):

you'll get an equation in terms of x and z

OpenStudy (anonymous):

so my finale for that would be -x-z=4 correct?

OpenStudy (anonymous):

@Kanwar245

OpenStudy (anonymous):

-x - z = 5

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

no what is the next step?

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

also you can rewrite last equation as 2x + 4 + z + 5z = -38

OpenStudy (anonymous):

2x + 6z = -42 or x + 2z = 31

OpenStudy (anonymous):

and you have -x -z = 5

OpenStudy (anonymous):

solve these two equations

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

you get z = 26

OpenStudy (anonymous):

x = -31

OpenStudy (anonymous):

is y 30?

OpenStudy (anonymous):

@Kanwar245

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Assuming Kanwar245 is correct, z = 26 x = -31 to find y, use these values in one of these equations to find y: x+y=-1 y-z=4 2x+y+5z=-38 from the first, x+y=-1, so \[\large y = -1 - x = -1 - (-31) = 30 \]. or use the second, y-z=4, so \[\large y= 4+z = 4+26 = 30\] or use the third, 2x+y+5z=-38, so \[y = -38 -2x -5z = -38 -2(-31)-5(26) = -106\] This doesn't agree with the other value for y of 30, so it's possible there's a mistake somewhere.